A local men's clothing store is being sold. The buyers are trying to estimate the percentage of items that are outdated. They will choose a random sample from the store's inventory in order to determine the proportion of merchandise that is outdated. We want our 95% confidence interval to have a margin of error of no more than plus or minus 3%. How large of a sample do the buyers need?
Answer #1

Solution :

Given that,

= 0.5

1 - = 0.5

margin of error = E = 0.03

Z/2 = 1.96

sample size = n = (Z / 2 / E)2 * * (1 - )

= (1.96 / 0.03)2 * 0.5 * 0.5

= 1067.1

sample size = n = 1068